OpenStudy (anonymous):

A system of equations is shown below. 3x + 8y = 12 2x + 2y = 3 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

help please ill give a medal

OpenStudy (the_fizicx99):

I'll use elimination method. 3x + 8y = 12 2x + 2y = 3  -6x -16y = -24 6x + 6y = 9 -10y = -15 /-10 /-10 y = 1.5 Sub in y into the equation and solve for x